OpenStudy (anonymous):

H-E-L-P Determine the extrema of below on the given interval f(x)=5x^3-61x^2+16x+3 (a) on [0,4] The minimum is and the maximum is . (b) on [-9,9] The minimum is and the maximum is

OpenStudy (anonymous):

first derivative and set to zero to find the local max and min other points of interest would be the end points

OpenStudy (anonymous):

THE DERIVATIVE IS 15X^2-122X+16

OpenStudy (anonymous):

DO WE FIND ALSO 2ND DERIVATIVE TOO?

OpenStudy (anonymous):

lol..here we are again...it's back to factoring.

OpenStudy (anonymous):

You need to find factors of 240, which is the product of 15 and 16 which add to the coefficient of the middle term.

OpenStudy (anonymous):

If you split the -122x into -120x and -2x, you will get the numbers I was referring to. Then you can factor by grouping.

OpenStudy (anonymous):

yup here we are... where the 240 come from...

OpenStudy (anonymous):

i dont believe 2nd derivative is necessary for determining extremas

OpenStudy (anonymous):

I explained already...240 comes from multiplying the 15(coefficient of squared term) and 16 (the constant).

OpenStudy (anonymous):

ok show it.

OpenStudy (anonymous):

why do you multiply it?

OpenStudy (anonymous):

and the middle term is -122x

OpenStudy (anonymous):

Because it is not a trinomial with leading coefficient of 1.

OpenStudy (anonymous):

yes. right..

OpenStudy (anonymous):

That is why I said you want numbers which add to the coeffcient of the middle term.

OpenStudy (anonymous):

The two numbers which multiply to give you 240 that add to -122 are -120 and -2.

OpenStudy (anonymous):

So, basic algebra tells us that if you want to factor an equation of the form \[ax ^{2}+bx+c=0 \] we must multiply a*c and find factors of a*c which add to the coeffcient of the middle term, in this case -122.

OpenStudy (anonymous):

And those factors, as I already mentioned, are -120 and -2. So, just attach the x to each of them, group the first two and last two terms in separate parenthesis and factor by grouping.

OpenStudy (anonymous):

\[15x ^{2}-122x+16=15x ^{2}-120x-2x+16=0\] Grouping them as I have instructed, you get \[(15x ^{2}-120x) + (-2x+16)=0\]

OpenStudy (anonymous):

Now just factor a 15x from the first parenthesis, and a -2 from the second parenthesis, and what do you notices when you do that?

OpenStudy (anonymous):

oh i got it... you just separated them ..

OpenStudy (anonymous):

can you separate them by anything? ex: -61x ?

OpenStudy (anonymous):

since 122/2 = 61..

OpenStudy (anonymous):

nvm i understand.. you want something you can easily factor with 15 and 2

OpenStudy (anonymous):

so i get.. x= 8 and 2/15

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

which one is max and min?

OpenStudy (anonymous):

wait the intervals are [0,4] what do we do with it?

OpenStudy (anonymous):

Check to see what happens to the derivative to the left and right of the point x=2/15 since x=8 is not in the interval.

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

true.. but what do you mean derivative to the left and right?

OpenStudy (anonymous):

plug x=0 and x=4 into the original equation in order to find the extrema points

OpenStudy (anonymous):

You can tell whether you have a max or min by what happens to the derivative to the left and right of the critical point. If f'(x) > o, your function is increasing, if f'(x)<0, your function is decreasing.

OpenStudy (anonymous):